# remarkRemark

### v1

@@ 5 Generalized training based on the PMP
(bSQHω) $\hat{\epsilon}_{k+1}=\begin{cases}\zeta\hat{\epsilon}_{k}&\text{if $j=0$}\\ \frac{1}{w+1}\sum_{i=k-\omega}^{k}\epsilon_{i}&\text{else }\end{cases},$
where $\zeta\in(0,1]$ and $\omega\in{\mathbb{N}}$, can be chosen suitably. The basis for this modification is the assumption that the choice of $\epsilon_{k}$ at each iteration implicitly depends on the local Lipchitz constants of the current mini-batch HP function. Thus, by taking the moving average of the approximations $\epsilon_{k}$ for these unknown constants, one expects to achieve an estimate for an optimal augmentation parameter in the next iteration. Even tough we do not give a rigorous proof for the effectiveness of our adjustment (bSQHω), we provide numerical evidence in Section 8 which shows that this adjustment can significantly reduce the number of forward steps required to determine a suitable augmentation parameter $\epsilon_{k}$ at each iteration. Moreover, while a suitable choice of the augmentation parameter can reduce the number of forward steps, introducing sparsity into the parameters of a CNN can reduce both the inference times while training and also the memory storage of a trained model.
@@ 6 Sparse network training based on the bSQH scheme
The most direct path to introduce sparsity into the solutions to (4) is a penalty term in the form the $L^{0}$-’norm’ which, for a given vector $v\in{\mathbb{R}}^{n}$, is defined by
@@
and thus strongly penalizes the number of non-zero parameters in a CNN. However, while this regularization technique is the most direct to introduce sparsity, it is seldom used in machine learning, due to its non-differentiability and optimization complexity, which yields classical backpropagation infeasible. The formulation of the updating step in the bSQH approach on the contrary, allows a straightforward derivation of an explicit update even in this non-continuous case. To prove this conjecture, we consider non-residual network structures such as those given in Proposition 3.1. Further, we assume the existence of a linear output layer, which can always be assumed without loss of generality by incorporating the non-linearity of the last layer into the terminal loss function. The corresponding forward propagation of such a network can be described as follows
@@
An important observation for our approach is that for a given set of parameters ${\mathbf{u}}$, both the network defined by (22) and the one defined corresponding to (23) have the same output $x_{L}$. Thus, for architectures such as (22), we can formulate (4) equivalently in terms of layer functions defined by (23). Due to the control-affine structure of (23), this in turn gives rise to the following augmented Hamiltonian function
${\cal H}^{B_{k}}_{l,\epsilon}({\mathbf{x}},{\mathbf{p}},w,u)={\cal F}^{B_{k}}_{l}({\mathbf{x}},{\mathbf{p}})u-\left(R_{l}(u)+\frac{\epsilon}{2}\|w-u\|^{2}\right),\quad{\cal F}^{B_{k}}_{l}({\mathbf{x}},{\mathbf{p}}):=\sum_{s\in I_{B_{k}}}p^{s}\cdot(F_{l}(x^{s})+{\cal B}_{l}),$
where $F_{l}(x^{s})$, for a fixed state $x^{s}$, denotes a suitable transformation matrix of the linear transformations $u\mapsto{\cal W}_{l}(u)f_{l-1}(x^{s})$ and $u\mapsto{\cal W}_{0}(u)x^{s}$ respectively. This formulation of the augmented Hamiltonian at hand, the updating-step in the bSQH algorithm simplifies to the following regularized quadratic problem
@@
$=\operatorname*{arg\,min}_{u\in U_{l}}\frac{1}{2}\left\|u-\left(w^{(k)}_{l}+\epsilon_{k}^{-1}{\cal F}^{B_{k}}_{l}({\mathbf{x}}^{(k)}_{l},{\mathbf{p}}^{(k)}_{l+1})\right)\right\|^{2}+\frac{\rho}{\epsilon_{k}}R_{l}(u).$
With a suitable regularizing function $R_{l}$, the optimization problem (24) admits a unique and analytic solution. Based on this reformulation and to enforce sparsity of a trained network by simultaneously increasing the stability of the forward propagation [7], we propose a regularization approach similar to the well-known Elastic-Net regularization [15], but with $L^{1}$-norm regularization replaced with (21) resulting in
@@
for all $l=0,\ldots,L-1$, where $\alpha\in[0,1)$. Due to the formulation of the updates in the form of (24), the trainable parameters at each iteration of Algorithm 1 take an analytic form for all $l=0,\ldots,L-1$, as follows
@@
$={\cal T}_{\frac{(1-\alpha)\rho}{\epsilon_{k}+\alpha\rho}}\left(\frac{\epsilon}{\epsilon_{k}+\alpha\rho}\left(u^{(k)}_{l}+\frac{1}{\epsilon_{k}}{\cal F}^{B_{k}}_{l}({\mathbf{x}}^{(k)}_{l},{\mathbf{p}}^{(k)}_{l+1})\right)\right),$
where we defined the Hard-threshold operator for a given vector $v\in{\mathbb{R}}^{n}$ by
${\cal T}_{\frac{(1-\alpha)\rho}{\epsilon_{k}+\alpha\rho}}(v)_{i}=\begin{cases}0&|v_{i}|\leq\sqrt{2\frac{(1-\alpha)\rho}{\epsilon_{k}+\alpha\rho}}\\ v_{i}&\text{otherwise}\end{cases},\qquad i=1,\ldots,n.$
Notice that the Elastic-Net approach can be recovered similarly by replacing the operator ${\cal T}_{(1-\alpha)\rho}$ in (26) with
${\cal S}_{\frac{(1-\alpha)\rho}{\epsilon_{k}+\alpha\rho}}(v)_{i}=\begin{cases}0&|v_{i}|\leq\frac{(1-\alpha)\rho}{\epsilon_{k}+\alpha\rho}\\ \left(v_{i}-\frac{(1-\alpha)\rho}{\epsilon_{k}+\alpha\rho}\right)\operatorname*{sgn}(v_{i})&\text{otherwise}\end{cases},\qquad i=1,\ldots,n.$
However, while such a penalty only encourages sparsity of the trained network parameters, a penalty of the form (25) enforces exact sparsity of a trained CNN. We present a comparison of the sparsity induced by both of these approaches in Section 8.
@@ 7 Convergence of the bSQH scheme
In the following, we provide results which prove that each updating step of the bSQH algorithm is feasible and can be generated within a finite number of steps. We remark that while some of these upcoming results are similar to those stated in [13], we present here a rigorous proof for the validity of the sufficient decrease condition
(27) $J_{B_{k}}\left({\mathbf{u}}^{(k+1)}\right)-J_{B_{k}}\left({\mathbf{u}}^{(k)}\right)\leq-\eta{|\kern-1.07639pt|\kern-1.07639pt|{\mathbf{u}}^{(k+1)}-{\mathbf{u}}^{(k)}|\kern-1.07639pt|\kern-1.07639pt|},$
under the relaxed assumptions A1 - A2, which include those discussed in [13] as a special case. As basis for the derivation of these results, we give the following auxiliary lemma.
@@
Let A1- A2 hold and define ${\mathbf{u}}\mapsto{\mathbf{x}}_{l}({\mathbf{u}})$ and ${\mathbf{u}}\mapsto{\mathbf{p}}_{l}({\mathbf{u}})$ for each ${\mathbf{u}}\in U_{ad}$ recursively by (8) and (9) respectively. Then for all $l=0,\ldots,L-1$. the control to adjoint map ${\mathbf{p}}_{l}(\;\cdot\;)$ is uniformly bounded and there exists a constant $C>0$ such that for any ${\mathbf{u}},{\mathbf{v}}\in U_{ad}$ it holds
(28) $\|{\mathbf{x}}_{l}({\mathbf{u}})-{\mathbf{x}}_{l}({\mathbf{v}})\|\leq C{|\kern-1.07639pt|\kern-1.07639pt|{\mathbf{u}}-{\mathbf{v}}|\kern-1.07639pt|\kern-1.07639pt|},\qquad\|{\mathbf{p}}_{l}({\mathbf{u}})-{\mathbf{p}}_{l}({\mathbf{v}})\|\leq C{|\kern-1.07639pt|\kern-1.07639pt|{\mathbf{u}}-{\mathbf{v}}|\kern-1.07639pt|\kern-1.07639pt|},$
@@
Due to the Lipschitz properties of $\Lambda_{l}$ for all $l=0,\ldots,L-1$, there exists a constant $C>0$ such that
(29) $\|{\mathbf{x}}_{l+1}({\mathbf{w}})-{\mathbf{x}}_{l+1}({\mathbf{u}})\|\leq C\left(\|{\mathbf{x}}_{l}({\mathbf{w}})-{\mathbf{x}}_{l}({\mathbf{u}})\|+\|w_{l}-u_{l}\|\right).$
@@
$\|{\mathbf{x}}_{l}({\mathbf{w}})-{\mathbf{x}}_{l}({\mathbf{u}})\|\leq C\sum_{j=0}^{l-1}\|w_{j}-u_{j}\|,$
which proves the Lipschitz property (28) of the control-to-feature maps. For the adjoint system, a similar reasoning applies. To prove the boundedness of the backpropagation variables, we have by the boundedness of $\partial_{x}\Lambda_{l}$, that there exists some constant $C>0$ for all $l=0,\ldots,L-1$ and any ${\mathbf{u}}\in U_{ad}$, such that
@@
By a change of index and an application of the discrete Gronwall Lemma, together with the Lipschitz continuity of $\Phi_{B}$, we have
$\|{\mathbf{p}}_{l}({\mathbf{u}})\|\leq C^{L}\|{\mathbf{p}}_{L}({\mathbf{u}})\|=C^{L}\|\nabla\Phi_{B}({\mathbf{x}}_{L}({\mathbf{u}}))\|\leq C^{L+1},$
for some generic constant $C>0$ and all ${\mathbf{u}}\in U_{ad}$, which gives the desired result.
@@
(31) $J_{B_{k}}({\mathbf{w}})-J_{B_{k}}({\mathbf{u}})\leq(C-\epsilon_{k}){|\kern-1.07639pt|\kern-1.07639pt|{\mathbf{w}}-{\mathbf{u}}|\kern-1.07639pt|\kern-1.07639pt|}.$
@@
For ease of notation assume $|B_{k}|=1$ and suppress the parameters $s$ and $B_{k}$. Remark that this adjustment can be made without loss of generality, due to the summable structure of the HP and cost functions. By definition, for any ${\mathbf{w}},{\mathbf{u}}\in U_{ad}$, we have:
@@
where $\Delta{\cal H}({\mathbf{w}},{\mathbf{u}}):=\sum_{l=0}^{L-1}{\cal H}_{l}({\mathbf{x}}_{l}({{\mathbf{u}}}),{\mathbf{p}}_{l+1}({\mathbf{u}}),w_{l})-{\cal H}_{l}({\mathbf{x}}_{l}({{\mathbf{u}}}),{\mathbf{p}}_{l+1}({\mathbf{u}}),u_{l})$. For the following, we define for all $l=0,\ldots,L-1$, the convex combination $x_{l}(\tau):={\mathbf{x}}_{l}({\mathbf{w}})+\tau({\mathbf{x}}_{l}({\mathbf{u}})-{\mathbf{x}}_{l}({\mathbf{w}}))$, where $\tau\in(0,1)$. Due to the convexity of ${\mathbb{X}}^{B}_{l}$ for all $l=0,\ldots,L-1$, the mean-value theorem guarantees the existence of $\tau_{1}\in(0,1)$ such that
@@
$=\left\langle\nabla_{x}{\cal H}_{l}(x_{l}(\tau_{1}),{\mathbf{p}}_{l+1}({\mathbf{u}}),w_{l}),{\mathbf{x}}_{l}({\mathbf{u}})-{\mathbf{x}}_{l}({\mathbf{w}})\right\rangle$
$=\langle{\mathbf{p}}_{l+1}({\mathbf{u}})^{T}\left(\partial_{x}\Lambda_{l}(x_{l}(\tau_{1}),w_{l})-\partial_{x}\Lambda_{l}({\mathbf{x}}_{l}({\mathbf{u}}),u_{l})\right),\left({\mathbf{x}}_{l}({\mathbf{u}})-{\mathbf{x}}_{l}({\mathbf{w}})\right)\rangle$
@@
Similarly, we have that there exists $\tau_{2}\in(0,1)$ such that
(34) $\Phi({\mathbf{x}}_{L}({\mathbf{w}}))-\Phi({\mathbf{x}}_{L}({\mathbf{u}}))=\langle\nabla\Phi(x_{L}(\tau_{2})),{\mathbf{x}}_{L}({\mathbf{u}})-{\mathbf{x}}_{L}({\mathbf{w}})\rangle.$
@@
$\quad+\langle\nabla\Phi(x_{L}(\tau_{2}))-\nabla\Phi({\mathbf{x}}_{L}({\mathbf{u}})),{\mathbf{x}}_{L}({\mathbf{w}})-{\mathbf{x}}_{L}({\mathbf{u}})\rangle$
$\quad+\sum_{l=0}^{L-1}\langle({\mathbf{p}}_{l+1}({\mathbf{u}})^{T}\left(\partial_{x}\Lambda_{l}(x_{l}(\tau),w_{l})-\partial_{x}\Lambda_{l}({\mathbf{x}}_{l}({\mathbf{u}}),u_{l})\right),{\mathbf{x}}_{l}({\mathbf{w}})-{\mathbf{x}}_{l}({\mathbf{u}})\rangle.$
@@
(36) $\|\nabla\Phi(x_{L}(\tau_{2}))-\nabla\Phi({\mathbf{x}}_{L}({\mathbf{u}}))\|\|{\mathbf{x}}_{L}({\mathbf{w}})-{\mathbf{x}}_{L}({\mathbf{u}})\|\leq(C+\tau_{2})\|x_{L}({\mathbf{w}})-{\mathbf{x}}_{L}({\mathbf{u}}))\|^{2}.$
@@
(37) $\langle({\mathbf{p}}_{l+1}({\mathbf{u}})^{T}\left(\partial_{x}\Lambda_{l}(x_{l}(\tau),w_{l})-\partial_{x}\Lambda_{l}({\mathbf{x}}_{l}({\mathbf{u}}),u_{l})\right),{\mathbf{x}}_{l}({\mathbf{w}})-{\mathbf{x}}_{l}({\mathbf{u}})\rangle$
$\qquad\leq\|{\mathbf{p}}_{l+1}({\mathbf{u}})\|\|\partial_{x}\Lambda_{l}(x_{l}(\tau),w_{l})-\partial_{x}\Lambda_{l}({\mathbf{x}}_{l}({\mathbf{u}}),u_{l})\|\|\left({\mathbf{x}}_{l}({\mathbf{w}})-{\mathbf{x}}_{l}({\mathbf{u}})\right)\|$
$\qquad\leq\|{\mathbf{p}}_{l+1}({\mathbf{u}})\|\left(C\|w_{l}-u_{l}\|\|{\mathbf{x}}_{l}({\mathbf{w}})-{\mathbf{x}}_{l}({\mathbf{u}})\|+(C+\tau_{1})\|{\mathbf{x}}_{l}({\mathbf{w}})-{\mathbf{x}}_{l}({\mathbf{u}})\|^{2}\right)$
Substituting the estimates determined in Lemma 7.1 into (35) and applying the Cauchy-Schwarz inequality, yields the existence of a constant $C>0$ such that for all ${\mathbf{w}},{\mathbf{u}}\in U_{ad}$ we have
(38) $J({\mathbf{w}})-J({\mathbf{u}})\leq-\Delta{\cal H}({\mathbf{w}},{\mathbf{u}})+C{|\kern-1.07639pt|\kern-1.07639pt|{\mathbf{w}}-{\mathbf{u}}|\kern-1.07639pt|\kern-1.07639pt|}.$
@@
(39) $\Delta{\cal H}({\mathbf{w}},{\mathbf{u}})\leq\epsilon_{k}{|\kern-1.07639pt|\kern-1.07639pt|{\mathbf{w}}-{\mathbf{u}}|\kern-1.07639pt|\kern-1.07639pt|}.$
@@
As a result of Theorem 7.3 the sufficient decrease condition (27) on the batch individual loss is guaranteed as soon as an augmentation parameter it determined satisfying $\epsilon_{k}>\eta+C$. Notice that this is ensured by the successive increase of the augmentation parameter $\epsilon_{k}$ corresponding to (20). In the particular case of the full-batch setting of Algorithm 1, this leads to a monotonically decreasing sequence of cost functional values.
@@
Suppose $B_{k}=B$ for all $k\in{\mathbb{N}}$ and let the sequence $({\mathbf{u}}^{(k)})_{k\in{\mathbb{N}}}\subset U_{ad}$ be defined such that $J_{B}({\mathbf{u}}^{(0)})<\infty$ and the sufficient decrease condition (27) holds for all $k\in{\mathbb{N}}$. Then if $J_{B}$ is bounded from below by zero and lower semicontinuous, the sequence $(J_{B}({\mathbf{u}}^{(k)}))_{k\in\mathbb{N}}$ of corresponding objective functional values is monotonically decreasing and $({\mathbf{u}}^{(k)})_{k\in{\mathbb{N}}}$ has a unique accumulation point $\bar{{\mathbf{u}}}\in U_{ad}$ with $J_{B}(\bar{{\mathbf{u}}})\geq\inf_{u\in U_{ad}}J_{B}({\mathbf{u}})$.

### v2

@@ 5 Generalized training based on the PMP
(bSQHω) $\hat{\epsilon}_{k+1}=\begin{cases}\zeta\hat{\epsilon}_{k}&\text{if }j=0\\ \frac{1}{\min(k+1,\omega+1)}\sum_{i=k-\min(k,\omega)}^{k}\epsilon_{i}&\text{else}\end{cases},$
where $\zeta\in(0,1]$ and $\omega\in{\mathbb{N}}$, must be chosen a-priori. The basis for this modification is the assumption that the choice of $\epsilon_{k}$ at each iteration implicitly depends on the local Lipschitz constants of the current mini-batch HP function. Thus, by taking the moving average of the approximations $\epsilon_{k}$ for these unknown constants, one expects to achieve an estimate for an optimal augmentation parameter in the next iteration. Even tough we do not give a rigorous proof for the effectiveness of our adjustment (bSQHω), we provide numerical evidence in Section 8 which shows that this adjustment can significantly reduce the number of forward steps required to determine a suitable augmentation parameter $\epsilon_{k}$ at each iteration. Moreover, while a suitable choice of the augmentation parameter can reduce the number of forward steps, introducing sparsity into the parameters of a CNN can reduce both the inference times and the memory storage of a trained model.
@@ 6 Sparse network training based on the bSQH scheme
The most direct path to introduce sparsity into the solutions to (4) is a penalty term in the form of the $L^{0}$-’norm’ which, for a given vector $v\in{\mathbb{R}}^{n}$, is defined by
@@
and thus strongly penalizes the number of non-zero parameters in a CNN. However, while this regularization technique is the most direct to introduce sparsity, it is seldom used in machine learning, due to its non-differentiability and optimization complexity, which makes classical backpropagation infeasible. The formulation of the updating step in the bSQH approach on the contrary, allows a straightforward derivation of an explicit update even in this non-continuous case. To prove this conjecture, we consider non-residual network structures similar to the well-known LeNet-5 architecture, where a linear output layer is used. We remark that a linear output layer can always be assumed without loss of generality by incorporating the non-linearity of the last layer into the terminal loss function. With the notation of Proposition 3.1, the corresponding forward propagation of such a network can be described as follows
@@
An important observation for our approach is that for a given set of parameters $\mathbf{u}$}, both the network defined by (21) and the one defined corresponding to (22) have the same output $x_{L}$. Thus, for architectures such as (21), we can formulate the supervised learning task (4) equivalently with neural networks constraints that admit layer architectures given by (22). Due to the control-affine structure of (22), this in turn gives rise to the following augmented Hamiltonian function
${\cal H}^{B_{k}}_{l,\epsilon}({\mathbf{x}},{\mathbf{p}},w,u)={\cal F}^{B_{k}}_{l}({\mathbf{x}},{\mathbf{p}})u-\left(\rho R_{l}(u)+\frac{\epsilon}{2}\|w-u\|^{2}\right),\quad{\cal F}^{B_{k}}_{l}({\mathbf{x}},{\mathbf{p}}):=\sum_{s\in I_{B_{k}}}p^{s}\cdot(F_{l}(x^{s})+{\cal B}_{l}),$
where $F_{l}(x^{s})$, for a fixed state $x^{s}$, denotes a suitable transformation matrix of the linear transformations $u\mapsto{\cal W}_{l}(u)f_{l-1}(x^{s})$ and $u\mapsto{\cal W}_{0}(u)x^{s}$ respectively. This formulation of the augmented Hamiltonian at hand, the updating step in the bSQH algorithm, simplifies to the following regularized quadratic problem
@@
$=\operatorname*{arg\,min}_{u\in U_{l}}\frac{1}{2}\left\|u-\left(w^{(k)}_{l}+\frac{1}{\epsilon_{k}}{\cal F}^{B_{k}}_{l}({\mathbf{x}}^{(k)}_{l},{\mathbf{p}}^{(k)}_{l+1})\right)\right\|^{2}+\frac{\rho}{\epsilon_{k}}R_{l}(u).$
With a suitable regularizing function $R_{l}$, the optimization problem (23) admits a unique and analytic solution. Based on this reformulation and to enforce sparsity of a trained network by simultaneously increasing the stability of the forward propagation [19], we propose a regularization approach similar to the well-known Elastic-Net regularization [45]. Specifically, we replace the $L^{1}$-norm in the Elastic-Net approach with (20) to obtain the following regularizer
@@
for all $l=0,\ldots,L-1$, where $\alpha\in[0,1)$. Due to the formulation (23), each update of the trainable parameters at each iteration of Algorithm 1 takes for all $l=0,\ldots,L-1$ the following explicit form
@@
$={\cal T}_{\frac{(1-\alpha)\rho}{\epsilon_{k}+\alpha\rho}}\left(\frac{\epsilon_{k}}{\epsilon_{k}+\alpha\rho}\left(u^{(k)}_{l}+\frac{1}{\epsilon_{k}}{\cal F}^{B_{k}}_{l}({\mathbf{x}}^{(k)}_{l},{\mathbf{p}}^{(k)}_{l+1})\right)\right)$,
where the Hard-threshold operator for a given vector $v\in{\mathbb{R}}^{n}$ and $\gamma>0$ is defined by
${\cal T}_{\gamma}(v)_{i}=\begin{cases}0&|v_{i}|\leq\sqrt{2\gamma}\\ v_{i}&\text{otherwise}\end{cases},\qquad i=1,\ldots,n.$
Notice that a similar reasoning applies in the Elastic-Net approach, where the $\cal T$} in (25) is replaced with the Soft-threshold operator defined as follows for a given vector $v\in{\mathbb{R}}^{n}$ and $\gamma>0$:
${\cal S}_{\gamma}(v)_{i}=\begin{cases}0&|v_{i}|\leq\gamma\\ \left(v_{i}-\gamma\right)\operatorname*{sgn}(v_{i})&\text{otherwise}\end{cases},\qquad i=1,\ldots,n.$
However, while the latter penalty only encourages sparsity of the trained network’s parameters, a penalty of the form (24) enforces exact sparsity of a trained CNN. We present a comparison of the sparsity-inducing properties of both approaches in Section 8.
@@ 7 Convergence of the bSQH scheme
In the following, we present results which prove that each updating step of the bSQH algorithm is feasible and can be generated within a finite number of steps. We remark that while some of these upcoming results are similar to those stated in [22], we present here a rigorous proof for the validity of the sufficient decrease condition
[…]
under the relaxed assumptions A1 - A2, which include those discussed in [22] as a special case. As a basis for the derivation of these results, we give the following auxiliary lemma.
@@
Let A1 - A2 hold and define ${\mathbf{u}}\mapsto{\mathbf{x}}_{l}({\mathbf{u}})$ and ${\mathbf{u}}\mapsto{\mathbf{p}}_{l}({\mathbf{u}})$ for each ${\mathbf{u}}\in U_{ad}$ recursively by (8) and (9) respectively. Then for all $l=0,\ldots,L-1$. the control to adjoint map ${\mathbf{p}}_{l}(\;\cdot\;)$ is uniformly bounded and there exists a constant $C>0$ such that for any ${\mathbf{u}},{\mathbf{w}}\in U_{ad}$ it holds
(27) $\|{\mathbf{x}}_{l}({\mathbf{u}})-{\mathbf{x}}_{l}({\mathbf{w}})\|\leq C\sum_{j=0}^{L-1}\|u_{j}-w_{j}\|,\qquad\|{\mathbf{p}}_{l}({\mathbf{u}})-{\mathbf{p}}_{l}({\mathbf{w}})\|\leq C\sum_{j=0}^{L-1}\|u_{j}-w_{j}\|,$
@@
Let ${\mathbf{u}},{\mathbf{w}}\in U_{ad}$, then due to the Lipschitz continuity of $\Lambda_{l}$ for all $l\in\{0,\ldots,L-1\}$, by assumption A2, there exists a constant $C>0$ such that
(28) $\|{\mathbf{x}}_{l+1}({\mathbf{u}})-{\mathbf{x}}_{l+1}({\mathbf{w}})\|\leq C\left(\|{\mathbf{x}}_{l}({\mathbf{u}})-{\mathbf{x}}_{l}({\mathbf{w}})\|+\|u_{l}-w_{l}\|\right).$
@@
$\|{\mathbf{x}}_{l}({\mathbf{u}})-{\mathbf{x}}_{l}({\mathbf{w}})\|\leq C\sum_{j=0}^{l-1}\|u_{j}-w_{j}\|,$
which proves the Lipschitz property (27) of the control-to-feature maps. For the adjoint map, a similar reasoning applies. To prove the boundedness of the backpropagation variables, we have by the boundedness of $\partial_{x}\Lambda_{l}$, that there exists some constant $C>0$ for all $l=0,\ldots,L-1$ and any ${\mathbf{u}}\in U_{ad}$, such that
@@
Thus, by a change of index and an application of the discrete Gronwall lemma, there exists a constant $C>0$ such that the following holds for all ${\mathbf{u}}\in U_{ad}$
$\|{\mathbf{p}}_{l}({\mathbf{u}})\|\leq C^{L}\|{\mathbf{p}}_{L}({\mathbf{u}})\|=C^{L}\|\nabla\Phi_{B}({\mathbf{x}}_{L}({\mathbf{u}}))\|.$
The boundedness of $\nabla\Phi_{B}$ by assumption A1 now proves the uniform boundedness of the parameter to adjoint maps.
@@
[…]
@@
For ease of notation, assume $|B_{k}|=1$ and suppress the parameters $s$ and $B_{k}$. It is worth noting that this adjustment can be made without loss of generality, due to the summable structure of the HP and cost function. By definition, for any ${\mathbf{w}},{\mathbf{u}}\in U_{ad}$, we have:
@@
where $\Delta{\cal H}({\mathbf{w}},{\mathbf{u}}):=\sum_{l=0}^{L-1}{\cal H}_{l}({\mathbf{x}}_{l}({{\mathbf{u}}}),{\mathbf{p}}_{l+1}({\mathbf{u}}),w_{l})-{\cal H}_{l}({\mathbf{x}}_{l}({{\mathbf{u}}}),{\mathbf{p}}_{l+1}({\mathbf{u}}),u_{l})$. For the following, we define for all $l=0,\ldots,L-1$, the convex combination $x_{l}(\tau)={\mathbf{x}}_{l}({\mathbf{w}})+\tau({\mathbf{x}}_{l}({\mathbf{u}})-{\mathbf{x}}_{l}({\mathbf{w}}))$, where $\tau\in(0,1)$. Due to the convexity of ${\mathbb{X}}^{B}_{l}$ for each $l=0,\ldots,L-1$, the mean-value theorem guarantees the existence of $\tau_{l}\in(0,1)$ such that
@@
$=\left\langle\nabla_{x}{\cal H}_{l}(x_{l}(\tau_{l}),{\mathbf{p}}_{l+1}({\mathbf{u}}),w_{l}),{\mathbf{x}}_{l}({\mathbf{u}})-{\mathbf{x}}_{l}({\mathbf{w}})\right\rangle$
$=\langle{\mathbf{p}}_{l+1}({\mathbf{u}})^{T}\left(\partial_{x}\Lambda_{l}(x_{l}(\tau_{l}),w_{l})-\partial_{x}\Lambda_{l}({\mathbf{x}}_{l}({\mathbf{u}}),u_{l})\right),\left({\mathbf{x}}_{l}({\mathbf{u}})-{\mathbf{x}}_{l}({\mathbf{w}})\right)\rangle$
@@
Similarly, we have that there exists $\tilde{\tau}\in(0,1)$ such that
(33) $\Phi({\mathbf{x}}_{L}({\mathbf{w}}))-\Phi({\mathbf{x}}_{L}({\mathbf{u}}))=\langle\nabla\Phi(x_{L}(\tilde{\tau})),{\mathbf{x}}_{L}({\mathbf{u}})-{\mathbf{x}}_{L}({\mathbf{w}})\rangle.$
@@
$\quad+\langle\nabla\Phi(x_{L}(\tilde{\tau}))-\nabla\Phi({\mathbf{x}}_{L}({\mathbf{u}})),{\mathbf{x}}_{L}({\mathbf{w}})-{\mathbf{x}}_{L}({\mathbf{u}})\rangle$
$\quad+\sum_{l=0}^{L-1}\langle({\mathbf{p}}_{l+1}({\mathbf{u}})^{T}\left(\partial_{x}\Lambda_{l}(x_{l}(\tau_{l}),w_{l})-\partial_{x}\Lambda_{l}({\mathbf{x}}_{l}({\mathbf{u}}),u_{l})\right),{\mathbf{x}}_{l}({\mathbf{w}})-{\mathbf{x}}_{l}({\mathbf{u}})\rangle.$
@@
(35) $\langle\nabla\Phi(x_{L}(\tilde{\tau}))-\nabla\Phi({\mathbf{x}}_{L}({\mathbf{u}})),{\mathbf{x}}_{L}({\mathbf{w}})-{\mathbf{x}}_{L}({\mathbf{u}})\rangle\leq(C+\tilde{\tau})\|{\mathbf{x}}_{L}({\mathbf{w}})-{\mathbf{x}}_{L}({\mathbf{u}}))\|^{2}.$
@@
(36) $\langle({\mathbf{p}}_{l+1}({\mathbf{u}})^{T}\left(\partial_{x}\Lambda_{l}(x_{l}(\tau_{l}),w_{l})-\partial_{x}\Lambda_{l}({\mathbf{x}}_{l}({\mathbf{u}}),u_{l})\right),{\mathbf{x}}_{l}({\mathbf{w}})-{\mathbf{x}}_{l}({\mathbf{u}})\rangle$
$\qquad\leq\|{\mathbf{p}}_{l+1}({\mathbf{u}})\|\|\partial_{x}\Lambda_{l}(x_{l}(\tau_{l}),w_{l})-\partial_{x}\Lambda_{l}({\mathbf{x}}_{l}({\mathbf{u}}),u_{l})\|\|\left({\mathbf{x}}_{l}({\mathbf{w}})-{\mathbf{x}}_{l}({\mathbf{u}})\right)\|$
$\qquad\leq\|{\mathbf{p}}_{l+1}({\mathbf{u}})\|\left(C\|w_{l}-u_{l}\|\|{\mathbf{x}}_{l}({\mathbf{w}})-{\mathbf{x}}_{l}({\mathbf{u}})\|+(C+\tau_{l})\|{\mathbf{x}}_{l}({\mathbf{w}})-{\mathbf{x}}_{l}({\mathbf{u}})\|^{2}\right)$
Substituting the estimates of (35) and (36) into (34) and using Lemma 7.1 together with the Cauchy-Schwarz inequality, guarantees the existence of a constant $C>0$ such that for all ${\mathbf{w}},{\mathbf{u}}\in U_{ad}$ we have
[…]
@@
[…]
@@
As a result of Theorem 7.3, the sufficient decrease condition (26) on the batch individual loss is guaranteed as soon as an augmentation parameter is determined satisfying $\epsilon_{k}>\eta+C$. Notice that this is ensured by the successive increase of the augmentation parameter $\epsilon_{k}$ corresponding to (19). In the particular case of the full-batch setting of Algorithm 1, this leads to a monotonically decreasing sequence of cost functional values.
@@
[…]